[time-nuts] HP 10811A failure

Adrian rfnuts at arcor.de
Sat Oct 13 19:26:35 EDT 2012


A very good point!

I checked the output amplitude with a spectrum analyzer, a power meter 
and a scope, the latter with a 50 ohm load to the input.
I re-checked the power supply connections and can confirm they are the 
same as before.
I used three different power supplies for the oscillators and two for 
the heaters, still no difference.
But...

To make it short, your comments helped to get me back on track. Both 
beauties escaped unnecessary dismantling and are now working as they should.

Actually, the problem was caused by a long-term misunderstanding. I was 
always wondering why 'HP used simple stranded wire' and not coaxial 
cable on the 10 MHz output and EFC.
As long as I connected them on the bench, I replaced the 'missing' 
ground connection with a short wired croc clamp between coax and ground. 
Today I realized that the thin blue wire IS actually coax cable, so I 
should have connected the coax shield to the BNC ground on my newly 
built 2x 10811 enclosure. With the new wiring, the output ground had 
just become much more inductive, up to a point where the nominal source 
impedance of 50 ohms had increased to over 120 ohms, causing the 
amplitude loss of some 3 dB and capturing noise.

Btw. the Sprague 6800 uF / 40 V from my R&S NGA power supply has indeed 
died, but independently of the osc. problem.

Adrian


GandalfG8 at aol.com schrieb:
> Perhaps a silly question, but as I get the impression that both seem  to
> have failed simultaneously with the same fault I'm just wondering  if you're
> sure it isn't what you're checking them with that's  developed a problem?
>   
> Regards
>   
> Nigel
> GM8PZR
>   
>   
> In a message dated 13/10/2012 13:52:46 GMT Daylight Time, rfnuts at arcor.de
> writes:
>
> Hi  All,
>
> both of my double oven 10811A's have been running flawlessly until  now
> when I noticed that the output power has dropped by about 3 dB  (measures
> only 4...4.5 dBm at 50 Ohm load), and the noise floor has gone  up by
> about 20 dB, while the frequency is still spot on.
>
> I used  them for various mesurements, so I can unfortunately not remember
> what  might have caused the failure. Only the 12V linear lab power supply
> that  fed the heaters had developed a completely dead filter cap, so the
> feeding  voltage was a 100 Hz sawtooth rather than DC. The heaters appear
> to be  still working though. I'm only using the internal heater. They
> draw some  700 mA at power up and drop to around 150 mA  each when warm.
> So the  oscillators are unlikely to be cooked inside. Something must have
> happened  to the 10 MHz output, possibly an ESD issue?
>
> Before I start taking them  apart, can anyone tell what has happened or
> what to look for  first?
>
> Regards,
> Adrian
>
>
>
> _______________________________________________
> time-nuts  mailing list -- time-nuts at febo.com
> To unsubscribe, go to
> https://www.febo.com/cgi-bin/mailman/listinfo/time-nuts
> and follow the  instructions there.
>
> _______________________________________________
> time-nuts mailing list -- time-nuts at febo.com
> To unsubscribe, go to https://www.febo.com/cgi-bin/mailman/listinfo/time-nuts
> and follow the instructions there.
>






More information about the time-nuts mailing list